# Training LLMs for Generating IEC 61131-3 Structured Text with Online Feedback

### v1

@@ Abstract
The advent of large language models (LLMs), such as GPT-3.5 and GPT-4, has enabled significant advancements in generating code across various domains. However, these models face unique challenges when generating IEC 61131-3 Structured Text (ST) code due to limited data in public training datasets and the complexity of ST language syntax. This paper proposes a novel approach to training LLMs that emphasizes improving the quality of learning data through an online process involving compiler feedback and evaluation from a secondary LLM. In this framework, the primary LLM generates new training samples, which are subsequently evaluated by a compiler for syntactical correctness and by a specialized LLM that excels at assessing semantic accuracy, though it is not optimized for code generation itself. Through iterative refinement of the training data, this approach results in marked improvements for the trained LLM, leading to higher compilation success rates and better semantic precision. As a result, the framework proves highly suitable for industrial automation applications and outperforms state-of-the-art models.
@@
IEC 61131-3; LLMs;
@@ I Introduction
We are entering a new era of code intelligence, where AI-driven copilots are transforming the way developers interact with code. These tools, powered by Large Language Models (LLM), streamline complex programming tasks and significantly enhance productivity [5]. However, as we approach a potential depletion of high-quality internet data by 2026, the reliance on synthetic data for model training will increase [21]. This shift presents significant challenges, particularly for niche programming languages such as IEC 61131-3 Structured Text (ST), which is prevalent in Programmable Logic Controller (PLC) engineering [2]. To ensure LLMs provide value in such specialized domains, it is essential that these models are aligned with human expectations and values, optimizing their outputs for real-world needs [6, 22].
@@
The first method introduced for model alignment, Reinforcement Learning from Human Feedback (RLHF) [9], trains a reward model (RM) using pairwise human preferences and then optimizes the policy through reinforcement learning (RL) based on the RM. Recently, Direct Alignment from Preferences methods (DAP) have emerged as popular alternatives to RLHF [24], [11]. Unlike RLHF, DAP methods update the language model (or policy) directly using pairwise preference data, making the alignment process more straightforward. However, the preference datasets are typically gathered prior to training and relying solely on human feedback to train these models can be prohibitively expensive, especially as LLMs grow in size and complexity. This is especially true for code, which requires domain experts due to its complexity. Therefore, innovative methods of alignment, beyond direct human feedback, will be necessary to maximize the benefits of these models.
In response to this challenge, this work proposes an iterative approach utilizing both compiler feedback and a dedicated semantic expert as evaluators. To this end and to overcome the data scarcity imposed by the PLC programming domain, the proposed framework exclusively exploits preexisting user prompts or intents, which consist of ST code descriptions and specifications, eliminating the need for large, labeled datasets. By iteratively refining the dataset and incorporating real-time feedback on both syntactic and semantic correctness, the approach ensures progressively better outcomes. This focus on improving the quality of the data, instead of relying solely on model improvements, leads to more reliable and semantically accurate code generation. The ultimate goal is to create a robust and scalable solution for industrial automation, overcoming the limitations of traditional human feedback and static datasets.
@@
II Existing State-of-the-Art Methods
Before giving details on the proposed methodology, let us take a quick look at a few related approaches.
@@ II Existing State-of-the-Art Methods
Initial approaches to LLM-based code generation involve using direct prompts, where models generate code from problem statements and example inputs [7]. More recent techniques, such as the chain-of-thought method [27], focus on breaking down tasks into modular or pseudo-code steps to improve planning and minimize errors. Furthermore, retrieval-based strategies enhance code generation by incorporating similar problems and their solutions to guide the process [23]. While these objectives are effective for natural language tasks, they fall short in fully capturing the complex syntactic and semantic rules of programming languages, often resulting in issues like syntax errors and type mismatches. Additionally, the success of these models in code generation is hindered by the need for large domain-specific datasets, which are typically limited in specialized fields like industrial PLCs. As a result, LLM-generated code often fails to pass test cases and lacks built-in mechanisms for automatic detection and correction of bugs [25].
Generative Pre-trained Transformer (GPT) models show great potential to generate PLC code based on domain-specific commands [18]. However, a significant challenge in this area remains the limited availability of data, which complicates effective model training and hinders performance. To address these challenges, the LLM4PLC project fine-tuned an open source model using publicly available PLC libraries such as OSCAT, demonstrating significant improvements in PLC code generation [12]. Despite recent advancements, models continue to struggle with generating fully accurate outputs on the first attempt, often requiring multiple corrections to resolve compilation errors. This highlights the ongoing need for refinement and optimization to achieve greater consistency with industrial standards. Code semantics can be effectively modeled through a dynamic environment that tracks variables in scope, including their types and other properties, as well as the functions being constructed. As the program progresses from start to finish, programming constructs alter this environment, influencing how subsequent code is interpreted. Therefore, the process of coding may be more accurately represented as a Markov Decision Process (MDP) rather than a simple next-token prediction task [22]. In this framework, the model receives positive rewards when it generates code that aligns with the given specifications. LLMs can be trained to succeed in this MDP using RL. This approach has already been applied to natural language generation, where human evaluators provide feedback in the form of rewards to guide the training process (RLHF) [26]. Recently, DAP methods have gained popularity as alternatives to RLHF.
@@
DAP methods, such as Direct Preference Optimization (DPO) and Kahneman-Tversky Optimization (KTO), update the target policy directly from preference pairs by training models with a preference model that simulates the evaluation of human feedback. In real-time, labelers compare model outputs to ensure consistency, but gathering large-scale annotations from human labelers is both expensive and logistically challenging. In most real-world applications, due to the high cost and complexity of collecting pairwise preferences, the preference dataset is typically compiled before aligning a language model and remains fixed throughout training. Gathering real-time preferences on new responses is typically impractical due to the absence of a human-in-the-loop for continuous feedback. As a result, the fixed data set makes all preference data static, preventing the policy from receiving feedback on its own generated output during the alignment process. Relevant works, such as Coarse-Tuning [17], CodeRL [19], StepCoder [10], and MapCoder [16], utilize compiler feedback to build preference datasets for model training. Another approach, Self-Play Fine-Tuning (SPIN) [8], employs an expert component, another instance of the generative model, to make the architecture self-sufficient. However, unlike preference-based methods, SPIN focuses on imitating a human-like style rather than leveraging explicit human preferences. Building on the concept of RL from AI Feedback (RLAIF), a novel approach called Online AI Feedback (OAIF) has been proposed for DAP methods [13]. This method combines the practical advantages of DAP with the online feedback mechanism of RLHF. Notably, OAIF bypasses the need for RM training by directly extracting preferences from the LLM.
III Proposed Method
The framework of the proposed method is based on the advantages of DAP techniques and real-time RLHF feedback, introducing compiler feedback to improve the syntactical correctness of the generated PLC code and unsupervised learning. By incorporating iterative refinement and domain-specific fine-tuning, we significantly enhance the accuracy and reliability of code generation, addressing gaps in prior approaches, and setting a potentially new standard for future research in this area. The flow chart representation of the proposed methodology is described in Fig. 1, and the detailed description of various blocks is as follows.
The key feature of this pipeline is its minimal data requirement, relying solely on intents $\mathcal{I}$, which are ST code descriptions and specifications, without the need for large, labeled datasets. The input set of intents, =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$, can be processed into smaller subsets $\mathcal{I}_{i}\subseteq\mathcal{I}$, where the size of each subset is determined by a sampling rate $\beta\in\mathopen{]}0,1\mathclose{]}$. The cardinality of each subset $\mathcal{I}_{i}$ is given by:
$Card(\mathcal{I}_{i})=\lfloor\beta\cdot Card(\mathcal{I})\rfloor.$ (1)
This allows for flexible sampling based on the chosen rate $\beta$, optimizing the data processing requirements for ST code generation.
Since in DPO, the off-policy issue persists during alignment as the policy gradually diverges from the supervised fine-tuning (SFT) baseline, and the RL step in RLHF is online with interactively acquired data, a straightforward solution is to introduce a reference model to provide continuous feedback. Using human feedback for this would be prohibitively expensive, but prior studies have shown that AI feedback is a reliable and effective for pairwise preference labeling [20]. Thus, we propose using an LLM as an online annotator to gather preference data over pairs of responses during alignment. This approach allows us to bypass traditional reward functions and policies by leveraging DPO, which enhances the model $\psi$ using syntactical and semantic preference assessments. Through iterative sampling, the model $\psi_{i}$ progressively improves, culminating in the optimized version $\psi_{DPO}$. Expert feedback plays a pivotal role in this refinement process, continuously improving the model’s comprehension and extending its applicability and scalability to a broad range of programming tasks — all achieved without the need for labeled data.
To improve the model’s generalization capabilities, the pipeline forgoes random sampling in favor of a more structured, deterministic approach that simulates batch processing. This method supports a stochastic fine-tuning mechanism, enabling more representative training iterations, which are essential for enhancing the model’s robustness and consistency.
In this context, let $\mathcal{\hat{C}}=\{\hat{C_{1}},\hat{C_{2}},\ldots,\hat{C_{n}}\}$ represent the set of synthetic code samples generated at each iteration. The model $\psi_{DPO}$ can be defined by the following function:
$\psi_{DPO}:\mathcal{I}\rightarrow\mathcal{\hat{C}}.$ (2)
At each iteration $i$, the model $\psi_{i-1}$ is employed to infer a subset $\mathcal{I}_{i}$, generating the corresponding synthetic set $\mathcal{\hat{C}}$. The baseline model $\psi$ thus serves as the initial LLM, denoted as $\psi_{0}$, for the pipeline.
The synthetic set $\mathcal{\hat{C}}$ is then divided into a positive code set $\mathcal{\hat{C}_{P}}$ and a negative code set $\mathcal{\hat{C}_{N}}$. The complete synthetic code set $\mathcal{\hat{C}}$ is defined as the union of these two sets:
$\mathcal{\hat{C}}=\mathcal{\hat{C}_{P}}\cup\mathcal{\hat{C}_{N}}.$ (3)
Prompting the annotating LLM to generate both positive and negative examples enables a structured online learning process. This approach continuously refines the model’s performance by using these examples to improve its learning. The resulting preference set is then defined as:
$\mathcal{C}_{DPO}=\mathcal{\hat{C}_{P}}\times\mathcal{\hat{C}_{N}}.$ (4)
The important step in creating the DPO dataset involves ranking the generated code samples into positive or negative categories, similar to the role of a reward function. This classification is performed by two distinct experts, each responsible for evaluating different aspects of the code.
The first expert, $\kappa$, is a compiler that provides dependable feedback on the syntactical correctness of the code. In addition to this syntactic check, a second expert is needed to assess the semantic accuracy, ensuring the code not only compiles but also correctly implements the intended logic. Rather than relying on extensive testing frameworks like unit tests, the LLM $\phi$ serves as the semantic expert, evaluating the correctness of the implementation without the need for external resources. Together, $\kappa$ and $\phi$ form a robust evaluation system for improving the model, ensuring both syntactical validity and semantic accuracy.
Given a set of synthetic code samples $\mathcal{\hat{C}}$ inferred from $\psi_{DPO}$, the functions $\mathcal{E}_{\kappa}$ and $\mathcal{E}_{\phi}$, which determine the syntactic and semantic validity of each code sample, are defined as follows:
$\begin{cases}\mathcal{E}_{\kappa}:\mathcal{\hat{C}}\rightarrow\{0,1\}\\[5.0pt] \mathcal{E}_{\phi}:\mathcal{\hat{C}}\rightarrow\{0,1\}\end{cases}$ (5)
In these functions, $\mathcal{E}(\hat{C_{i}})=1$ indicates that the code sample $\hat{C_{i}}$ is correct, while $\mathcal{E}(\hat{C_{i}})=0$ signifies either compilation errors (in the case of $\mathcal{E}_{\kappa}$) or semantic inconsistencies (in the case of $\mathcal{E}_{\phi}$).
Using the results of $\mathcal{E}_{\kappa}$ and $\mathcal{E}_{\phi}$, a ranking policy is established to sort the code samples within $\mathcal{\hat{C}}$. For any two code samples $\hat{C_{i}},\hat{C_{j}}\in\mathcal{\hat{C}}$ where $i,j\in\{1,\ldots,Card(\mathcal{\hat{C}})\}$, the following ranking rule is applied:
$\mathcal{E}_{\phi}(\hat{C_{i}})\leq\mathcal{E}_{\kappa}(\hat{C_{i}})<\mathcal{E}_{\phi}(\hat{C_{j}})\leq\mathcal{E}_{\kappa}(\hat{C_{j}})\iff\hat{C_{i}}\prec\hat{C_{j}}.$ (6)
This rule implies that the ranking of code samples is determined by their syntactic and semantic validity, with $\hat{C_{i}}$ ranked lower than $\hat{C_{j}}$ if it scores worse in both syntactical ($\mathcal{E}_{\kappa}$) and semantic ($\mathcal{E}_{\phi}$) evaluations. Given two code samples $\hat{C_{i}}$ and $\hat{C_{j}}$, the expert $\mathcal{E}_{\phi}$ is used in order to compare them and select the best one in terms of syntax and semantics in case $\hat{C_{i}}\sim\hat{C_{j}}$. The sample $C_{p}$ is then the unique positive sample, and all the elements contained in $\mathcal{C}\setminus\{C_{p}\}$ are considered as negative. This process enables the establishment of a reliable ground truth based on the set $\mathcal{C}$, while simultaneously constructing a dataset for DPO, denoted as $\mathcal{D}$. This dataset $\mathcal{D}$ serves as a foundation for training and refining the model, ensuring that both syntactic and semantic aspects are properly accounted for during the optimization process.
As the dataset evolves and becomes more refined, the model $\psi_{DPO}$ undergoes fine-tuning, closely following a traditional training loop but with specific adjustments to minimize the risk of overfitting. In each iteration, a distinct subset $\mathcal{D}_{i}\subseteq\mathcal{D}$, composed of both positive and negative code pairs, is chosen for further refinement. This strategic selection of data ensures that the model is exposed to a diverse set of examples, enabling better generalization capabilities as the dataset, not just the model, improves with each iteration.
In alignment with the sampling phase, the model $\psi_{i-1}$ undergoes fine-tuning at each iteration $i$. The iterative training process is structured as follows:
  • Use the dataset $\mathcal{D}$ to fine-tune $\psi_{i-1}$ via DPO, resulting in the updated model $\psi_{i}$. The most recent DPO model is then updated such that $\psi_{DPO}=\psi_{i}$.
  • Evaluate the performance of $\psi_{i}$ on a validation set to track improvements and adjust hyperparameters as needed for optimal performance.
To ensure the integrity of the model $\psi_{DPO}$ and prevent it from diverging too far from its initial output distribution, the baseline model $\psi$ is designated as the reference model throughout the DPO process. This reference model acts as an anchor, ensuring that as $\psi_{DPO}$ evolves through fine-tuning, it retains the core characteristics and performance benchmarks of $\psi$. Finally, the iterative process is characterized by:
$\theta_{i}=\arg\max_{\psi_{i-1}}\left(\mathbb{E}_{(I_{j},C_{j})\sim\mathcal{D}}\left[\log\frac{\psi_{i-1}(C_{j}\mid I_{j})}{\psi(C_{j}\mid I_{j})}\right]-\lambda\mathcal{R}(\theta_{i-1})\right),$ (7)
where $\theta$ represents the fine-tuning parameters of a model $\phi$.
Through this structured and iterative fine-tuning approach, the dataset used to train $\psi_{DPO}$ improves over time, becoming more refined with each iteration. This gradual enhancement of the data, facilitated by the preference optimization mechanism, allows the model to better learn new patterns while preserving its foundational knowledge. As the quality of the training data improves, $\psi_{DPO}$ achieves better generalization across various applications, ensuring a balance between acquiring new insights and maintaining consistency.

### v3

@@ Abstract
IEC 61131-3 Structured Text (ST) is a widely used programming language for programmable logic controllers (PLCs) in automation systems. However, generating ST code with LLMs poses unique challenges due to limited data in public training datasets and the complexity of ST language syntax. This paper proposes an approach to fine-tune LLMs for the generation of ST code that leverages a preference-based learning method through an online process involving compiler feedback and evaluation from an LLM-based ST expert. In this framework, the model is iteratively refined and generates new training samples, which are subsequently evaluated by a compiler for syntactical correctness and by a specialized LLM that excels at assessing semantic accuracy, though it is not optimized for code generation itself. This approach results in marked improvements for the trained LLM, leading to higher compilation success rates and better semantic precision. As a result, the framework proves highly suitable for industrial automation applications and outperforms state-of-the-art models.
@@
IEC 61131-3; LLMs; LLM Fine-tuning; Compiler feedback;
@@ I Introduction
We are entering a new era of code intelligence, where AI-driven copilots are transforming how developers interact with code. These tools, powered by Large Language Models (LLM), streamline complex programming tasks and significantly enhance productivity [1]. However, as we approach a potential depletion of high-quality internet data by 2026, the necessity of synthetic data for model training will increase [2]. This shift presents significant challenges, particularly for niche programming languages such as IEC 61131-3 Structured Text (ST), which is prevalent in Programmable Logic Controller (PLC) engineering and industrial automation systems [3]. To ensure that LLMs provide value in such specialized domains, it is essential that these models are aligned with expert expectations and values, optimizing their output for real-world needs [4, 5].
@@
Reinforcement Learning from Human Feedback (RLHF) [6], a pioneering LLM alignment method, trains a reward model (RM) using pairwise human preferences and then optimizes the policy through reinforcement learning (RL) based on the RM. Recently, Direct Alignment from Preferences methods (DAP) have emerged as popular alternatives to RLHF [7], [8]. Unlike RLHF, DAP methods directly update the language model using pairwise preference data, simplifying the alignment process because there is no need to train a dedicated reward model. In the original DAP approach [9], preference datasets are gathered prior to training and are based on expert feedback. With LLMs growing in size and complexity, this method can become prohibitively expensive if human experts label IEC 61131-3 conform generated code. Therefore, innovative alignment methods beyond direct human feedback are needed to fully leverage these models.
In response to this challenge, this work utilizes synthetic experts in an iterative online approach based on syntactic compiler feedback and a dedicated semantic LLM expert. To overcome the data scarcity of PLC code samples, the proposed framework exclusively exploits preexisting user prompt and code pairs derived from the APPS dataset [10] for SFT, as well as additional prompts that are subsequently used in an online finetuning scheme similar to Online AI Feedback (OAIF) [11]. This eliminates the need for large code datasets. The approach iteratively refines the model and incorporates feedback on both syntactic and semantic correctness, ensuring progressively better outcomes.
We base our proposed method on three concepts to overcome data scarcity in the ST code generation domain. First, there exists a performance margin between code generation and code quality judgement. Datasets used for training State-of-the-Art GPTs include only a small number of ST code samples, or none at all. Even though these models may not be able to generate ST code themselves, they are able to judge ST code samples due to semantic similarities with other languages like Pascal, which they were trained on extensively. Similarly to concepts like OAIF [11] or [12] we use the margin between generation and evaluation performance of SOTA LLMs to enhance ST code generation. While DPO typically applies to static offline preference use, recent studies found that DPO also benefits from iterative online training [13, 14, 15].
Second, using compiler feedback allows an accurate mapping from a diverse distribution of code samples to a binary compile label for each intent input [16]. This potentially enables training on an infinite amount of synthetically generated and syntactically flawless code samples beyond the syntax judgment performance of LLM judges.
Third, leaning on the concept of natural selection, a set of many code samples is generated instead of only code sample pairs. This enables the syntactic (compiler feedback) and semantic judge (LLM) to explore a broader distribution of samples [17]. It was shown that a high number of samples results in improved output quality [18, 17]. Hence, we believe that a sufficient number of code samples has to be generated to find a syntactically and semantically flawless code sample.
@@
II State-of-the-Art Methods
Before giving details on the proposed methodology, we investigate state-of-the-art methods for LLMs for code generation, OAIF and compiler feedback.
II-A LLM-based Code Generation
@@
Initial approaches for LLM-based code generation took problem statement prompts as input and directly generated code outputs conditioned on the prompts [19]. More recent techniques, such as chain-of-thought prompting [20] focus on breaking down tasks into modular or pseudo-code steps to improve planning and minimize errors. Furthermore, retrieval-based strategies incorporate similar problem samples and code solutions to guide the code generation process [21]. While these objectives are effective for natural language tasks, they fall short in fully capturing the complex syntactic and semantic rules of programming languages, often resulting in issues like syntax errors and type mismatches. Additionally, the success of these models in code generation is hindered by the need for large domain-specific datasets, which are scarce for specialized fields like industrial PLCs. As a result, LLM-generated code often fails to pass test cases and lacks built-in mechanisms for automatic detection and correction of syntactic and semantic errors [22].
II-B Recent advances in LLM Alignment
Generative Pre-trained Transformer (GPT) models show great potential to generate PLC code based on domain-specific commands [23]. However, a significant challenge in this area remains the limited availability of data, which complicates effective model training and hinders performance. To address these challenges, the LLM4PLC project fine-tuned an open source model using publicly available PLC libraries such as OSCAT, demonstrating significant improvements in PLC code generation [24]. Despite recent advances, models still struggle with first-attempt code accuracy, typically requiring multiple iterations to resolve compilation errors. For effective use of code in industrial applications, further refinements are needed.
@@ II-B Recent advances in LLM Alignment
An approach that greatly improves code quality is Reinforcement Learning for Human Feedback (RLHF). In RLHF, human evaluators provide preference data by comparing pairs of model outputs. This data is used to train a reward model (RM), which is incorporated into the LLM model training process [5]. Recently, DAP methods have gained popularity as alternatives to RLHF. DAP methods, such as Direct Preference Optimization (DPO) [9], update the target policy directly from preference pairs by training models with a preference model that simulates the evaluation of human feedback. In real-time, labelers compare model outputs to ensure consistency, but gathering large-scale annotations from human labelers is both expensive and logistically challenging. In most real-world applications, due to the high cost and complexity of collecting pairwise preferences, the preference dataset is typically compiled before aligning a language model and remains fixed throughout training. Gathering real-time preferences on new responses is typically impractical due to the absence of a humans-in-the-loop for continuous feedback. As a result, the fixed dataset makes all preference data static, preventing the policy from receiving feedback on its own generated output during the alignment process. Relevant works, such as Coarse-Tuning [25], CodeRL [26], StepCoder [27], and MapCoder [28] utilize compiler feedback to improve model training. Recent advances use DPO teratively, e.g. [9][13].
II-C LLM-based Structured Text Code Generation
Structured Text is one of the programming languages defined in the standard IEC 61131-3 for programmable logic controllers and has been designed for automation and control tasks. As the necessity of LLMs able to generate ST code is prevalent, several groups have developed methods for ST code generation, recently. Multiple base LLM models like GPT-3 and fine-tuned GPT-J have already been finetuned for generating ladder logic, function block diagrams, and ST code [29]. The models were found to generate basic PLC code, but struggled with more complex logic and integration with real-world PLC hardware and libraries. Another group showed that retrieval-augmented ST code generation integrates function blocks from libraries like OSCAT into the LLM-generated code [30]. LLM4PLC further introduced syntax checkers and model verifiers and utilized fine-tuning with Low-Rank Adaptations (LoRAs) to adapt LLMs for ST code generation. The applied methods showed a significant improvement in ST code generation performance, evaluated by domain experts [31]. The summarized papers were all published in 2024 and demonstrate the growing industrial interest for ST code generation. Despite these efforts, improvements are required to meet industry requirements and shrink the margin between finetuned models for ST code and SOTA models for frequently used programming languages like Python.
III Method
This section describes the two subsequent stages for fine-tuning the LLM with special emphasis on the online-dpo method.
III-A Supervised Fine-Tuning
Supervised Fine-Tuning (SFT) adapts a pre-trained language model (LLM) $\theta_{0}$ for domain-specific tasks. During SFT, the model is trained on a set of instruction-response pairs $D_{\text{sft}}=\{(d_{1},d^{r}_{1}),\dots,(d_{m},d^{r}_{m})\}$. This step ensures the model can follow instructions accurately and apply specific syntax, such as ST. SFT bridges the gap between general language understanding from pre-training and the specialized requirements of downstream tasks by maximizing the likelihood of generating the correct response $d^{r}$ (i.e. ST control code) given the instruction $d$ (i.e. code intent), as described in Equation 1.
$\mathcal{L}_{\text{SFT}}=\arg\max_{\theta_{0}}\sum_{(x,y)\in D_{\text{sft}}}\log\pi(y|x)$ (1)
SFT is conducted as the first step of fine-tuning, using an instruction-response dataset consisting of user intents and corresponding code. This assumes that a sufficient number of pairs is available. For ST, there are only a few public datasets (e.g., OSCAT [32]). The synthetic generation of such datasets can be time-consuming and expensive, especially if the model has limited understanding of code.
III-B Direct-Preference-Optimization with AI and Compiler-Feedback
DPO (Direct Preference Optimization) is a widely used technique for model alignment based on expert feedback. In contrast to SFT (Supervised Fine-Tuning), it utilizes an instruction and preference triple rather than an instruction-response pair. This means that instead of requiring a code example, it only needs to be determined which output from a model, based on a given prompt, is preferred over another. This approach has the advantage of exploiting formal evaluation methods, such as unit tests for code as an expert. An actuall reference code sample is not needed. Classical offline DPO uses static preference datasets and conducts a single training round on them [19]. These datasets can be generated in various ways, such as by querying a better model (a form of knowledge extraction) or by using the model being trained. If the model lacks sufficient performance, sampling methods (e.g., [smaller models source]) or self-correction approaches based on feedback could be employed [source]. However, this has the drawback of requiring a large number of samples to obtain high-quality positive preference pairs, or it may lead to a distribution shift between the preference samples and the actual model, which could result in degraded learning performance [source?]. To achieve continuous improvement of the model, we employ an iterative online variant of DPO [29][30]. This newer DPO training technique ensures that the preference pairs remain aligned with the model’s distribution, also allowing the model to benefit from its own improvements in generating positive samples [source]. In this approach, the model being aligned generates samples, which are then used as preference pairs for fine-tuning through DPO. This process is repeated iteratively: the updated model generates new outputs, which are again evaluated by the expert to create new preference samples for further fine-tuning. This has the advantage that only instructions and an expert for the preference collection are needed for training, rather than instruction-response pairs with code, which are often scarce in the case of SFT for code.
III-C Description of preference collection
The formal description of the procedure is as follows.
Let the instruction set be defined as: $\mathcal{X}\triangleq\{x_{1},x_{2},\dots,x_{M}\}$ where $M=|\mathcal{X}|$ denotes the cardinality of the set.
For each iteration $i\in\{1,\dots,I\}$, a subset $X_{i}\triangleq\{X_{i,1},\dots,X_{i,N}\}$ is drawn from $\mathcal{X}$: $X_{i}\sim\pi_{i}(\mathcal{X}),$ where $\pi_{i}(\mathbf{X})$ represents the selection strategy applied during iteration $i$ and $N$ is the number of drawn instructions. For each instruction, a number of $T$ code responses from the model $\theta_{i}$ is sampled, the set is denoted by $Y_{i}\triangleq\{Y_{i,1},\dots,Y_{i,T}\}$. In this paper, two experts are queried to build preferences over the responses. The first expert is the compiler, which checks the syntactic correctness. The second expert is an LLM, which is used as a soft substitute for unit tests. The experts are combined to generate the preference pairs in the following way. Each generated code sample in $Y_{i}$ is tested on the compilation. The samples that compile, are checked by the semantic expert using a LLM. The samples that do not compile, are treated as negative samples. The LLM as semantic expert then labels the compiling samples again as positive or negative, depending on their semantics. All positive samples are then combined with all negative samples. In case no samples are positive after the compile check, all samples are given to the LLM and labeled.
The combination of the respective expert feedback ultimately results in a preference dataset $D_{i}$, where $Y_{w}\succ Y_{l}$:
$D_{i}=\{(X_{i,n},y_{w,n},y_{l,n})\}\forall n\in\{1,\dots,N\},$
III-D Direct-Preference-Optimization Fine-Tuning
After collecting the prefence dataset $D_{i}$ in the $i$th iteration, a model training leverging DPO is conducted and the model parameters are updated to $\theta_{i+1}$. The overall process is depicted in Fig. 1.
Algorithm 1 Fine-Tuning algorithm to train a LLM for ST code generation
Input:
  • Number of training iterations $I$
  • Instruction dataset $\mathcal{X}$
  • Supervised dataset $\mathcal{D_{\text{sft}}}$
  • Code generation model $\theta_{0}$
  • Structured Text experts (semantic and syntax experts)
1:Initialize $\theta$ with supervised fine-tuning (SFT) to get $\theta_{0}$
2:for $i\leftarrow 1$ to $I$ do
3:     Sample $N$ ST intents $X_{i}\sim\pi_{i}(\mathcal{X})$
4:     Generate $T$ ST code responses from $\theta_{i-1}$ for each intent
5:     Leverage expert feedback to construct preference dataset $D_{i}$
6:     Update model $\theta_{i}$ using $D_{i}$ via DPO
7:end for
Output: Aligned code generation model $\theta_{I}$